# Data-driven model order reduction for wave propagation in materials with nonlinearities or damage

### v1

@@ 3.6 Dynamic mode decomposition
Dynamic mode decomposition (DMD) is a purely data-driven approach proposed in [Schmid2010] to extract spatio-temporal coherent structures from high-dimensional dynamical systems. For a comprehensive study of the method, we refer to [BN19, HTu2014, KBBP16], for example. The basic idea is to treat a sequence of measurement snapshots as observations of an underlying linear operator (often interpreted as a finite-dimensional approximation of the Koopman operator, see [ROWLEY2009]). In order to characterize the system’s dynamics, DMD computes eigenvalues and eigenvectors of this linear operator. The resulting modes evolve exponentially in time allowing to identify growth rates or oscillatory patterns without explicitly knowing the underlying governing equations.
@@
It can be shown that the reduced operator $\tilde{\mathbf{A}}$ has the same eigenvalues as the full order operator $\mathbf{A}$, see [HTu2014]. We use the notation $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\dots,\lambda_{r})$ and refer to $(\lambda_{i})_{i=1}^{r}$ as the DMD eigenvalues. The eigenvectors of the reduced operator $\tilde{\mathbf{A}}$ can be utilized in order to construct the DMD modes [^2] Note that we use the notation $\mathbf{\Phi}$ to refer to the DMD modes in this Section and the POD modes in Section 3.3; the intended interpretation is clear from the context. $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$ as
@@
These DMD modes are the eigenvectors of the full order operator $\mathbf{A}$ such that we can compute the spectral decomposition of $\mathbf{A}$ without the need to carry out explicit computations on the matrix $\mathbf{A}$ directly. Utilizing the spectral decomposition $(\mathbf{\Phi},\mathbf{\Lambda})$ it is possible to reconstruct the system state, here written in a formulation which is continuous in time (see e.g. [BN19, (7.27)]):
@@
For the computation of the initial mode amplitudes $\mathbf{b}$, we consider two choices in this work: the first option (see e.g. [BN19, (7.29)]) is given by
@@
where the second expression computes the amplitudes using POD projected data. As second option, the optimal vector of amplitudes $\mathbf{b}$ could be computed by solving an optimization problem as mentioned in [Jovanovi2014] (section II.B). The final formula for the optimal $\mathbf{b}$ is:
@@
While classical DMD constructs spatio-temporal modes that are coherent over the entire time series, a well-known challenge arises in case phenomena appear only locally over short intervals. In order to overcome this limitation, multiresolution DMD (mrDMD) was introduced in [Kutz2016]. The main idea of mrDMD is to combine DMD with a hierarchical, time-localized analysis and thereby separating the dynamics across multiple temporal scales. This not only retains the long-lasting global behaviors captured by classical DMD but also identifies localized patterns. In practice, the application of mrDMD involves recursively partitioning the time domain into different time levels with smaller sub-windows and applying DMD at each level of resolution. We denote the time domain split at level $j$ and sub-window $w$ as: $\mathbf{X}_{j,w}=[\mathbf{x}_{w_{1}},\mathbf{x}_{w_{2}},\dots,\mathbf{x}_{w_{k}}]$, here $w_{1},w_{2},…,w_{k}$ represent the time indices in the sub-window. At each level $j$, for sub-window $w$, standard DMD is applied to $\mathbf{X}_{j,w}$, yielding:
@@ 4 Numerical examples
This section presents the results of applying the model order reduction techniques of Section 3 to three different numerical examples, some of which involve localized damage. The first example considers the wave equation (1) in a two-dimensional domain, where the damage is modeled as a localized reduction in the wave speed coefficient. For this benchmark problem, we assume full access to the governing equations and simulation data, enabling a comprehensive comparison of both intrusive and non-intrusive MOR techniques. The second example involves a complex linear elasticity model of a composite material, specifically fiber metal laminate (FML), in which the damage is represented by a localized reduction in the Young’s modulus. Here, we apply the data-driven, non-intrusive MOR methods DMD and OpInf, reflecting a practical scenario where knowledge about the structure of the system is available as well as input-output data. The third example concerns GUW propagation in a nonlinear elastic model of intact aluminium, without damage. Here, we focus exclusively on data-driven DMD assuming only displacement data is available and the full order model is not accessible.
@@ 4.1 Numerical results for the wave equation
In this subsection, we consider the problem of the wave equation (1) with a parameterized damage and recall that this model is taken from [Bur2024], where the problem setting is based on [NandaLorenzWave]. For the initial boundary value problem of the wave equation (1) we make the following choices: the spatial domain is $\Omega=(0,5)\times(0,5)\subset\mathbb{R}^{2}$ and the time interval is set to $(t_{0},t_{end})=(0.1,5)$. We set the right-hand side $f(t,\mathbf{x})=0$ and the boundary condition $b(t,\mathbf{x})=0$. The initial conditions are chosen as $u_{0}(\mathbf{x})=u_{\mathbf{x}_{M}}(0.1)g(\mathbf{x})$ and $u_{1}(\mathbf{x})=\frac{du_{\mathbf{x}_{M}}}{dt}(0.1)g(\mathbf{x})$, where $g$ is defined as follows:
@@
with $\mathbf{x}=(x_{1},x_{2})\in\Omega$ and $u_{\mathbf{x}_{M}}$ is given by
@@
where $\mathbf{x}_{D}=(x_{1D},x_{2D})$ is the location of the damage at which the value of the wave speed is reduced to $c_{D}<0.25$ and $d_{D}$ is a parameter which describes the size of the damage. Piecewise linear finite elements with $h_{max}=0.05$ are used to discretize (1) in space leading to $N=11317$ spatial degrees of freedom and the Crank-Nicolson method is used to solve the discretized system in time starting from $t_{0}=0.1$ with a time step value of $\tau=0.02$ which gives $M=245$ snapshots. Figure 1 shows the resulting numerical solution for the displacement of the finite element model at the times $t=0,2.5,5$ with the damage at position $x_{D}=(3,3)\in\Omega$ with $d_{D}=0.06$ and a reduction of the wave speed to $c_{D}=0.05$. The wave motion is excited around the center $(2.5,2.5)\in\Omega$ and propagates outwards from there. The excited wave is reflected at the damage in position $\mathbf{x}_{D}$ resulting in a second overlaid wave motion centering around the damage.
@@
Figure 1: The solution of the FOM for the wave equation at times $t=0,2.5,5.$
@@
In the following, we compare the solutions to the reduced order models constructed using POD, DMD and OpInf in terms of accuracy and efficiency. The implementation details are given as follows: The DMD initial amplitudes are computed using two options: (i) projection of the first snapshot onto the POD basis (see (46)), and (ii) an optimized amplitude formulation based on a least-squares fit over the time series (see (47)). We refer to these two versions of DMD by ’DMD’ and ’DMD Optimal’, respectively. We mention here that the DMD solution is obtained using (41). Regarding operator inference, we assume that we have full knowledge about the external forces which are assumed to be identically zero for all time instants. This corresponds to the forces-informed variant of OpInf. Along the solution data, the approach requires the data for the second order derivative of the displacement mentioned in (16). In this work, the required acceleration (second time derivative) data is computed using an eighth order finite difference scheme, see [Sharma2024], chosen to ensure high accuracy in the OpInf optimization problem. In order to improve solution stability during the optimization, a Tikhonov regularization term, scaled by the Frobenius norm of the inferred operators, is added to the OpInf optimization problem (29). The optimization problem solves for the auxiliary matrices $\mathbf{L}$ and $\mathbf{W}$ which are subsequently used to compute the reduced mass and stiffness matrices. The optimization is implemented using the TensorFlow library [TensorFlow], with the Adam algorithm [Adam] used as the optimizer. To account for scale differences between the displacement and its second derivative, the scaling method described in Section 3.5 is applied.
Figure 3 shows a comparison of the solution of the full order model to the solution of the reduced order model at the sensor location for the reduced dimensions $r=40$ and $r=80,$ respectively for POD, DMD, Optimal DMD and OpInf.
@@
We notice that the POD-ROM is able to reproduce qualitatively the full order model signal. As for the quantitative assessment of the POD results, it yields a relative sensor error $\epsilon_{s}$ (51) of $2.331$ $\%$ and $2.13\times 10^{-6}$ $\%$ for the reduced dimensions of $40$ and $80$, respectively. As for DMD and Optimal DMD, we observe that both versions have reconstructed the FOM solution successfully. The relative error values of DMD and Optimal DMD for the case of $r=40$ are $5.137$ $\%$ and $2.702$ $\%$, respectively. As for the case of $r=80$, the relative error values of both DMD versions are $0.231$ $\%$ and $0.003$ $\%$, respectively. For the OpInf-based ROM, the case $r=40$ resulted in a relatively high error of $25.03$ $\%$ with a clear amplitude offset in the second half of the dynamics. Increasing the reduced dimension to $r=80$ substantially enhanced the model accuracy, yielding an error of only $4.36$ $\%$.
@@
Next, we compare the decay of the relative error $\epsilon_{s}$ at the sensor location $(3.5,2.5)\in\Omega$ depending on the reduced dimension $r$ for the reduced order solutions generated by POD, DMD, Optimal DMD and OpInf in Figure 4 (left). For POD-ROM, we observe that the relative error $\epsilon_{s}$ decays rapidly and falls below $10\%$ for $r=21$ and below $5\%$ for $r=29$. DMD and Optimal DMD have similar convergence pattern with the latter being more accurate for all selected reduced dimensions, the Optimal DMD sensor error goes below the 33 $\%$ threshold at $r=40$. As for OpInf ROM, the sensor error for $r=20$ and $r=40$ is significantly high, then the quality of the approximation of OpInf becomes higher with $r=60$ and then starts to saturate at $r=80$ with error value of $1.824$ $\%$. The time-dependent relative $L^{2}$-error $\epsilon_{u}(t)$ (52) over the entire spatial domain for the ROM of dimension $r=40$ is plotted in Figure 4 (right). For POD-ROM and Optimal DMD, most of the simulation, the error stays between $3\%$ and $5\%$, while DMD gives slightly higher error values. In case of OpInf-ROM, the error value grows as time passes. This is attributed to the fact that the reduced inferred operators are not identical to the projected true ones which results in an error that accumulates throughout the simulation time. However, these results could be improved as will be mentioned later in this subsection.
@@
Figure 3: The full order displacement versus the reduced order approximation at the sensor location for the reduced dimensions $r=40$ (left) and $r=80$ (right) for the wave equation, respectively.
Figure 4: Relative $L^{2}$-error $\epsilon_{s}$ at the sensor location as a function of the reduced dimension $r$ (left) and relative $L^{2}$-error $\epsilon_{u}(t)$ over the entire spatial domain as a function of time for the reduced dimension $r=40$ (right) for reduced order solutions computed using POD, DMD and OpInf for the wave equation. The values are in percentages.
For the purpose of illustration, the displacement computed with the POD reduced order model of dimension $r=40$ (top) as well as the difference between the solutions to the FOM and the POD-ROM (bottom) at the times $t=0,2.5$ and $t=5$ are shown in Figure 5.
(a) $t=0$
(b) $t=2.5$
(c) $t=5$
Figure 5: The POD reduced order solution of dimension $r=40$ (top) and the difference between the full order and POD reduced order solution (bottom) for the wave equation at times $t=0,2.5$ and $t=5.$
Let us further study the effect of using more snapshot data in the operator inference approach for the accuracy of the reduced order model by reducing the time step size. The displacement time signals at the sensor location are plotted in Figure 6 (left) for both the full order solution and the reduced one based on OpInf using a time step of $0.02$ $\mathrm{s}$. The relative error $\epsilon_{s}$ is $7.162\%$. We improve the results using more snapshot data by utilizing a smaller time step of $0.01$ $\mathrm{s}$ which doubles the number of snapshots from $245$ to $490$. The OpInf procedure is repeated with the new dataset, and the reduced system is integrated using the new time step. As shown in Figure 6 (right), the alignment between the FOM and the ROM improves significantly, reducing the relative error $\epsilon_{s}$ to $2.885\%$.
Refer to captionRefer to caption
@@
Figure 6: The full order time signal of the displacement at the sensor location versus the reduced order one obtained by OpInf for $M=245$ (left) and $M=490$ (right), $r=80$ for the problem of the wave equation.
@@
Finally, we compare the reduced order methods in terms of computational complexity. Table 1 summarizes the offline and online times for POD, DMD and OpInf for different time steps sizes. We would like to mention that the offline times include the FOM simulation time which is $4.086$ $\mathrm{s}$ and $8.057$ $\mathrm{s}$ for the choices of the time steps of $0.02$ $\mathrm{s}$ and $0.01$ $\mathrm{s}$, respectively. Clearly, OpInf has very large offline computational times compared to POD and DMD. However, the online computational times are comparable for all methods, with DMD appearing to be the fastest.
@@
Table 1: The computational times of the different reduced order methodologies applied for the first numerical example of the wave equation with damage. The unit is $\mathrm{s}$].
Time Step ROM Type POD-ROM DMD OpInf-ROM
Offline Online Offline Online Offline Online
$\tau=0.02$ $5.193$ $0.023$ $4.163$ $0.015$ $672.486$ $0.034$
$\tau=0.01$ $9.883$ $0.049$ $8.289$ $0.031$ $1001.657$ $0.069$

### v2

@@ 3.8 Dynamic mode decomposition
Dynamic mode decomposition (DMD) is a purely data-driven approach proposed in [76] to extract spatio-temporal coherent structures from high-dimensional dynamical systems. For a comprehensive study of the method, we refer to [13, 36, 50], for example. The basic idea is to treat a sequence of measurement snapshots as observations of an underlying linear operator (often interpreted as a finite-dimensional approximation of the Koopman operator, see [74]). In order to characterize the system’s dynamics, DMD computes eigenvalues and eigenvectors of this linear operator. The resulting modes evolve exponentially in time allowing to identify growth rates or oscillatory patterns without explicitly knowing the underlying governing equations.
@@
It can be shown that the reduced operator $\tilde{\mathbf{A}}$ has the same non-zero eigenvalues as the full order operator $\mathbf{A}$, see [36]. We use the notation $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\dots,\lambda_{r})$ and refer to $(\lambda_{i})_{i=1}^{r}$ as the DMD eigenvalues. The eigenvectors of the reduced operator $\tilde{\mathbf{A}}$ can be utilized in order to construct the DMD modes [^2] Note that we use the notation $\mathbf{\Phi}$ to refer to the DMD modes in this Section and the POD modes in Section 3.2; the intended interpretation is clear from the context. $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$ as
@@
These DMD modes are the eigenvectors of the full order operator $\mathbf{A}$ such that we can compute the spectral decomposition of $\mathbf{A}$ without the need to carry out explicit computations on the matrix $\mathbf{A}$ directly. Utilizing the spectral decomposition $(\mathbf{\Phi},\mathbf{\Lambda})$ it is possible to reconstruct the system state, here written in a formulation which is continuous in time (see e.g. [13, (7.27)]):
@@
For the computation of the initial mode amplitudes $\mathbf{b}$, we consider two choices in this work: the first option (see e.g. [13, (7.29)]) is given by
@@
where the last expression computes the amplitudes using POD projected data. As second option, the optimal vector of amplitudes $\mathbf{b}$ could be computed by solving an optimization problem as mentioned in [47] (Section II.B). The final formula for the optimal $\mathbf{b}$ is:
@@
While classical DMD constructs spatio-temporal modes that are coherent over the entire time series, a well-known challenge arises in case phenomena appear only locally over short intervals. In order to overcome this limitation, multiresolution DMD (mrDMD) was introduced in [51]. The main idea of mrDMD is to combine DMD with a hierarchical, time-localized analysis and thereby separating the dynamics across multiple temporal scales. This not only retains the long-lasting global behaviors captured by classical DMD but also identifies localized patterns. In practice, the application of mrDMD involves recursively partitioning the time domain into different time levels with smaller sub-windows and applying DMD at each level of resolution. We denote the time domain split at level $j$ and sub-window $w$ as: $\mathbf{X}_{j,w}=[\mathbf{x}_{w_{1}},\mathbf{x}_{w_{2}},\dots,\mathbf{x}_{w_{k}}]$, here $w_{1},w_{2},…,w_{k}$ represent the time indices in the sub-window. At each level $j$, for sub-window $w$, standard DMD is applied to $\mathbf{X}_{j,w}$, yielding:
@@ 4 Numerical examples
This section presents the results of applying the model order reduction techniques of Section 3 to three different numerical examples, some of which involve localized damage. The first example considers the wave equation (1) in a two-dimensional domain, where the damage is modeled as a localized reduction in the wave speed coefficient. For this benchmark problem, we assume full access to the governing equations and simulation data, enabling a comprehensive comparison of both intrusive and non-intrusive MOR techniques. The second example involves a complex linear elasticity model of a composite material, specifically fiber metal laminate (FML), in which the damage is represented by a localized reduction in the Young’s modulus. Here, we apply the data-driven, non-intrusive MOR methods DMD and OpInf, reflecting a practical scenario where knowledge about the structure of the system is available as well as input-output data. The third example concerns GUW propagation in a nonlinear elastic model of aluminium. Here, we focus exclusively on data-driven DMD assuming only displacement data is available and the full order model is not accessible.
@@ 4.1 Numerical results for the wave equation
In this subsection, we consider the problem of the wave equation (1) with a parameterized damage and recall that this model is taken from [15], where the problem setting is based on [55]. For the initial boundary value problem of the wave equation (1) we make the following choices: the spatial domain is $\Omega=(0,5)\times(0,5)\subset\mathbb{R}^{2}$ and the time interval is set to $(t_{0},t_{end})=(0.1,5)$. We set the right-hand side $f(t,\mathbf{x})=0$ and the boundary condition $b(t,\mathbf{x})=0$. The initial conditions are chosen as $u_{0}(\mathbf{x})=u_{\mathbf{x}_{M}}(0.1)g(\mathbf{x})$ and $u_{1}(\mathbf{x})=\frac{du_{\mathbf{x}_{M}}}{dt}(0.1)g(\mathbf{x})$, where $g$ is defined as follows:
@@
with $\mathbf{x}=(x_{1},x_{2})\in\Omega,\quad\mathbf{x}_{c}=(x_{c},y_{c})=(2.5,2.5)\in\Omega$ and $u_{\mathbf{x}_{M}}$ is given by
@@
where $\mathbf{x}_{D}=(x_{1D},x_{2D})$ is the location of the damage at which the value of the wave speed is reduced to $c_{D}<0.25$ and $d_{D}$ is a parameter which describes the size of the damage. Piecewise linear finite elements with $h_{max}=0.05$ are used to discretize (1) in space leading to $N=11317$ spatial degrees of freedom and the Crank-Nicolson method is used to solve the discretized system in time starting from $t_{0}=0.1$ with a time step value of $\tau=0.02$ which gives $M=245$ snapshots. The initial fields of the displacement and its time derivative are shown in (1).
(a) $\mathbf{u}_{0}$
(b) $\dot{\mathbf{u}}_{0}$
Figure 1: Initial displacement and initial time derivative of the displacement fields for the problem of the wave equation
Figure 2 shows the resulting numerical solution for the displacement of the finite element model at the times $t=0,2.5,5$ with the damage at position $x_{D}=(3,3)\in\Omega$ with $d_{D}=0.06$ and a reduction of the wave speed to $c_{D}=0.05$. The wave motion is excited around the center $\mathbf{x}_{c}=(2.5,2.5)\in\Omega$ and propagates outwards from there. The excited wave is reflected at the damage in position $\mathbf{x}_{D}$ resulting in a second overlaid wave motion centering around the damage.
@@
Figure 2: Solution of the FOM for the wave equation at times $t=0,2.5,5.$
@@
In the following, we compare the solutions to the reduced order models constructed using POD, DMD and OpInf in terms of accuracy and efficiency. The implementation details are given as follows: The DMD initial amplitudes are computed using two options: (i) projection of the first snapshot onto the POD basis (see (61)), and (ii) an optimized amplitude formulation based on a least-squares fit over the time series (see (62)). We refer to these two versions of DMD by ’DMD’ and ’DMD Optimal’, respectively. We mention here that the DMD solution is obtained using (56).
@@
Regarding operator inference, we apply both continuous and discrete version of OpInf to this problem. We describe firstly the procedure for the time continuous setting. In this problem we have full knowledge about the external forces which are assumed to be identically zero for all time instants. This corresponds to the forces-informed variant of OpInf. Along the solution data, the approach requires the data for the second order derivative of the displacement mentioned in (20). The required acceleration (second time derivative) data is computed using an eighth order finite difference scheme, see [77], chosen to ensure high accuracy in the OpInf optimization problem. In order to improve solution stability during the optimization, a Tikhonov regularization term, scaled by the Frobenius norm of the inferred operators, is added to the OpInf optimization problem (26). The optimization problem solves for the auxiliary matrices $\mathbf{L}$ and $\mathbf{E}$ which are subsequently used to compute the reduced mass and stiffness matrices. The optimization is implemented using the TensorFlow library [1], with the Adam algorithm [48] used as the optimizer. To account for scale differences between the displacement and its second derivative, the scaling method described in Section 3.5 is applied. It has to be noted that in this example the zero operator $\mathbf{\widehat{P}}=\mathbf{0}$ is a trivial solution to the minimization problem because of the fact that the right hand side is identically zero. Therefore, a suitable initialization of the auxiliary matrices has to be done in order to avoid the trivial solution.
We shift the discussion to the time discrete version of OpInf which we denote by dOpInf. We aim at learning the reduced order operators $\mathbf{\widehat{H}}$ and $\mathbf{\widehat{R}}$ in (52). We remark that $\mathbf{\widehat{C}}=0$ because of the absence of external input signal or loads. In this numerical example we have applied dOpInf without and also with the use of the re-projection method in (3.7). For both cases of dOpInf, we have generated snapshots which correspond to $q=6$ different initial data. In particular we sampled uniformly $x_{c}$ in (69) in the range of $[2.5,3.5]$ and we fixed $y_{c}=2.5$. The generated snapshots are then concatenated to give the following matrices:
$\mathbf{U}_{0}=[\mathbf{U}_{0}^{1},\mathbf{U}_{0}^{2},\ldots,\mathbf{U}_{0}^{q}],\quad\mathbf{U}_{1}=[\mathbf{U}_{1}^{1},\mathbf{U}_{1}^{2},\ldots,\mathbf{U}_{1}^{q}],\quad\mathbf{U}_{2}=[\mathbf{U}_{2}^{1},\mathbf{U}_{2}^{2},\ldots,\mathbf{U}_{2}^{q}]].$ (72)
We then apply the re-projection method in Algorithm 2 to each initial data which gives the re-projected trajectories $\mathbf{\bar{U}}_{0}^{1},\dots\mathbf{\bar{U}}_{0}^{q}$, $\mathbf{\bar{U}}_{1}^{1},\dots\mathbf{\bar{U}}_{1}^{q}$ and $\mathbf{\bar{U}}_{2}^{1},\dots\mathbf{\bar{U}}_{2}^{q}$, concatenating the re-projected trajectories gives:
$\mathbf{\bar{U}}_{0}=[\mathbf{\bar{U}}_{0}^{1},\mathbf{\bar{U}}_{0}^{2},\ldots,\mathbf{\bar{U}}_{0}^{q}],\quad\mathbf{\bar{U}}_{1}=[\mathbf{\bar{U}}_{1}^{1},\mathbf{\bar{U}}_{1}^{2},\ldots,\mathbf{\bar{U}}_{1}^{q}],\quad\mathbf{\bar{U}}_{2}=[\mathbf{\bar{U}}_{2}^{1},\mathbf{\bar{U}}_{2}^{2},\ldots,\mathbf{\bar{U}}_{2}^{q}].$ (73)
Then the data matrices $\mathbf{\widehat{D}}$ and $\mathbf{\bar{D}}$ are assembled as in (52) and (54) for the dOpInf without and with the re-projection method, respectively. In this setting, the matrices $\mathbf{\widehat{D}}$ and $\mathbf{\bar{D}}$ are both full rank which implies that the minimization problems correspond to solving two sets of matrix equations. The accuracy of the learned operators depends highly on the condition number of the matrices $\mathbf{\widehat{D}}\mathbf{\widehat{D}}^{T}$ and $\mathbf{\bar{D}}\mathbf{\bar{D}}^{T}$. We solve the system of equations to obtain the learned operators $\mathbf{\widehat{P}}$ and $\bar{\mathbf{{P}}}$ by inverting the matrices $\mathbf{\widehat{D}}\mathbf{\widehat{D}}^{T}$ and $\mathbf{\bar{D}}\mathbf{\bar{D}}^{T}$, respectively.
These computations were carried out using Matlab 2023b.
@@
At this point, we present the results of the reduction performed by POD-ROM, DMD and OpInf. We start by visualizing the results obtained by POD-ROM and DMD versus the FOM. As an example, we consider the reduced dimensions of $20$ and $40$. The displacement solutions at the sensor location for POD, DMD and optimal DMD are plotted in Fig. 4. We notice that the POD-ROM is able to reproduce qualitatively the full order model signal. As for the quantitative assessment of the POD results, it yields a relative sensor error (66) $\epsilon_{s}=8.872$ $\%$ and $3.021$ $\%$ for the reduced dimensions of $20$ and $40$, respectively. As for DMD and optimal DMD, we observe that both versions have reconstructed the FOM solution successfully. The relative error values of DMD and optimal DMD for the case of $r=20$ are $13.371$ $\%$ and $9.671$ $\%$, respectively. As for the case of $r=40$, the relative error values of both DMD versions are $5.137$ $\%$ and $2.702$ $\%$, respectively. The results here indicate a more accurate reconstruction of the FOM signal is obtained via the optimal DMD variant with respect to the classical one. In addition we notice that optimal DMD and POD-ROM are close to each other in terms of accuracy.
We now present the results obtained using operator inference. In particular, we compare the performance of time continuous and time discrete operator inference against the POD-ROM and the FOM. For consistency, the same displacement signal shown in Fig. 4 is considered for all approaches, and the corresponding comparisons are shown in Fig. 5. The results indicate that the time continuous operator inference approach is not sufficiently accurate in reconstructing the full order model (FOM) signal, as a clear deviation is observed during the second half of the time history for both reduced dimensions $r=20$ and $r=40$. In contrast, the final plot shows that the time discrete operator inference approach achieves a more accurate quantitative reconstruction of the FOM signal, both with and without the use of re-projection.
To provide a quantitative assessment, time continuous OpInf yields relative sensor errors $\epsilon_{s}=16.812$ $\%$ and $24.943$ $\%$ for $r=20$ and $r=40$, respectively. For time discrete OpInf without re-projection, the corresponding sensor errors are $16.884$ $\%$ and $4.832$ $\%$. Finally, time discrete OpInf with re-projection achieves the highest accuracy among all tested reduced order models, with sensor errors of $8.532$ $\%$ and $2.4$ $\%$.
These results demonstrate that time discrete operator inference with re-projection is a highly accurate reduction methodology for second order mechanical systems. The improved performance relative to the case without re-projection can be attributed to a more accurate recovery of POD-intrusive operators enabled by the re-projection procedure. In fact this can be seen by computing the relative error between the learned operator in case dOpInf with and without re-projection as follows:
$\epsilon_{H}=\frac{\left\lVert\operatorname{vec}(\mathbf{H}_{r})-\operatorname{vec}(\mathbf{H}^{\star})\right\rVert_{L^{2}(0,T)}}{\left\lVert\operatorname{vec}(\mathbf{H}_{r})\right\rVert_{L^{2}(0,T)}}\times 100\%,\quad\epsilon_{R}=\frac{\left\lVert\operatorname{vec}(\mathbf{R}_{r})-\operatorname{vec}(\mathbf{R}^{\star})\right\rVert_{L^{2}(0,T)}}{\left\lVert\operatorname{vec}(\mathbf{R}_{r})\right\rVert_{L^{2}(0,T)}}\times 100\%,$ (74)
where here $\operatorname{vec}$ is the vectorization operator and the operators $\mathbf{H}^{\star}$ and $\mathbf{R}^{\star}$ are the learned operators of dOpInf either $\mathbf{\bar{H}},\,\mathbf{\bar{R}}$ or $\widehat{\mathbf{H}},\,\widehat{\mathbf{R}}$. In the current example, the maximum value of $\epsilon_{H}$ for dOpInf with re-projection is $1.201$ $\%$ at $r=100$, while the corresponding $\epsilon_{R}$ remains below $1\times 10^{-7}$ $\%$ for all reduced dimensions. In contrast, for dOpInf without re-projection, the lowest value of $\epsilon_{H}$ is $3.956$ $\%$ at $r=20$, whereas the highest value reaches $25.118$ $\%$ at $r=90$. For the same model, the minimum value of $\epsilon_{R}$ is $7.779$ $\%$ at $r=30$, while the maximum value is $39.102$ $\%$ at $r=90$.
We plot the condition number of the matrices $\widehat{\mathbf{D}}\widehat{\mathbf{D}}^{T}$ and $\bar{\mathbf{D}}\bar{\mathbf{D}}^{T}$ versus the reduced dimension in Fig. 6. As it can be seen from the last plot, the condition number of the re-projection case ($\bar{\mathbf{D}}\bar{\mathbf{D}}^{T}$) is always smaller than the one without re-projection ($\widehat{\mathbf{D}}\widehat{\mathbf{D}}^{T}$).
@@
Next, we compare the decay of the relative error $\epsilon_{s}$ at the sensor location $(3.5,2.5)\in\Omega$ depending on the reduced dimension $r$ for the reduced order solutions generated by POD, DMD, Optimal DMD and OpInf in Fig. 7 (left). For POD-ROM, we observe that the relative error $\epsilon_{s}$ decays rapidly and falls below $10\%$ for $r=20$ and below $5\%$ for $r=30$. DMD and Optimal DMD have similar convergence pattern with the latter being more accurate for all selected reduced dimensions, the Optimal DMD sensor error goes below the $3\%$ threshold at $r=40$. As for OpInf ROM, the sensor error for $r=20$ and $r=40$ is significantly high, then the quality of the approximation of OpInf becomes higher with $r=60$ and then starts to saturate at $r=80$ with error value of $4.36$ $\%$. As for the dOpInf approaches, we notice that both perform better than their continuous counterpart. In particular, dOpInf without re-projection has achieved an error value of $2.4\%$ for $r=50$, however, dOpInf with re-projection has resulted in low error values even lower than the ones obtained by POD-ROM for some choices of $r$. For example, for $r=50$ the dOpInf with re-projection has yielded an error of $1.84\%$ (lower than the POD-ROM with $2.322\%$) and it has given values lower than $1\%$ for $r\geq 60$.
The time-dependent relative $L^{2}$-error $\epsilon_{u}(t)$ (67) over the entire spatial domain for the ROM of dimension $r=40$ is plotted in Fig. 8. For POD-ROM and Optimal DMD, most of the simulation the error stays between $3\%$ and $5\%$, while DMD gives slightly higher error values. In case of OpInf, the error value grows as time passes. This is attributed to the fact that the reduced inferred operators are not identical to the projected true ones which results in an error that accumulates throughout the simulation time. On the other hand, more accurate results are obtained using the discrete variants (dOpInf), where the one without re-projection is giving better results (than OpInf) for the first $4$ seconds of the simulation. The dOpInf with re-projection is the most accurate reduction methodology in this test with an error value even slightly better than the intrusive POD-ROM at certain time values.
@@
Figure 4: Full order displacement versus POD-ROM and DMD solutions at the sensor location for the reduced dimensions $r=20$ (left) and $r=40$ (right) for the wave equation, respectively.
@@
Figure 5: Full order displacement versus POD-ROM, OpInf and dOpInf solutions at the sensor location for the reduced dimensions $r=20$ (left) and $r=40$ (right) for the wave equation, respectively.
Figure 6: Condition number for the matrices $\widehat{\mathbf{D}}\widehat{\mathbf{D}}^{T}$ and $\bar{\mathbf{D}}\bar{\mathbf{D}}^{T}$ which are inverted to solve the dOpInf problem without and with re-projection, respectively.
Figure 7: Relative $L^{2}$-error $\epsilon_{s}$ at the sensor location as a function of the reduced dimension $r$ for POD-ROM compared with DMD variants (left) and POD-ROM compared with variants of OpInf (right).
Figure 8: Relative $L^{2}$-error $\epsilon_{u}(t)$ over the entire spatial domain as a function of time for the reduced dimension $r=40$ for reduced order solutions computed using POD, DMD variants (left), and OpInf and dOpInf variants (right) for the wave equation.
@@
Finally, we compare the reduced order methods in terms of computational complexity. Table 1 summarizes the offline and online times for POD, DMD, OpInf and dOpInf variants for the reduced dimension of $r=40$. We would like to mention that the offline times include the FOM simulation time which is $4.086$ $\mathrm{s}$. Clearly, OpInf has very large offline computational times compared to POD and DMD. However, the online computational times are comparable for all methods. The offline time needed for dOpInf with re-projection is obviously larger than dOpInf without re-projection because of the additional offline time required to query the FOM in the re-projection method.
@@
Table 1: Computational times of different reduced order methodologies for the wave equation with damage for $r=40$. The unit is $\mathrm{s}$].
Stage ROM Type POD-ROM DMD OpInf dOpInf dOpInf Re-projection
Offline 5.193 4.163 672.486 27.655 76.351
Online 0.0149 0.013 0.0156 0.0108 0.0120